OpenStudy (anonymous):

Hunus! n-5/2<_2n-6/3

OpenStudy (rogue):

\[(n-5)/2 \le (2n-6)/3\]Cross multiply.\[3(n-5) \le 2(2n-6)\]\[3n - 15 \le 4n - 12\]Subtract 3n from both sides, and add 12 to both sides.\[3n - 15 - 3n + 12 \le 4n - 12 - 3n + 12\]\[-3 \le n\]